Morgan Stanley has a current cash flow (at time 0) of $3.4 m and pays no dividends. The present value of the company’s future cash flows is $14.6 m. The firm is entirely financed with equity and has 400,000 shares outstanding. Assume the dividend tax rate is zero.




1. Suppose the board of directors of Morgan Stanley announces its plan to payout 40% of its current cash flow as dividends to its shareholders. How can Andy, who owns 800 shares of Morgan Stanley stock, achieve a zero payout policy on his own?






1
Expert's answer
2022-02-17T11:06:38-0500

First, compute the stock price.


Stock Price = "\frac{PV of all Cash Flows}{Total Shares Outstanding}"

Present Value of all cash flows = Current Period's Cash Flow + Present Value of Future Cash Flows

Current period's cash flow = $3.4m

Present Value of future cash flows = $14.6m

Present Value of all cash flows = 3.4m + 14.6m

= $18m

Total Shares Outstanding = 400,000

Stock Price = "\frac{18,000,000}{400,000}"

= $45


Then, find the total dividends Andy will receive

Dividend Paid Out = Payout Rate "\times" Cash Flow

If Payout Rate = 40%, & Cash Flow = $3.4m

Dividend Paid Out = 40% "\times" $3.4

= $1,360,000


Andy will receive:

=> Total dividend paid out "\times" "\frac{Andy's Shares}{Total Number of Shares}"

=> $1,360,000 "\times" "\frac{800}{400,000}"

= $2720


Then, find the new stock price, for a zero payout policy to hold true.

New Stock Price = Old Stock Price - "\frac{Total Dividend Paid Out}{Total Number of Shares}"

= $45 - "\frac{1,360,000}{400,000}"

= $41.6


To achieve a zero payout policy, Andy needs to purchase a number of stocks as shown below,

Additional Stocks to Purchase = "\frac{Total Dividend Received by Andy}{New Share Price}"

= "\frac{2720}{41.6}"

= 65.38 (2 decimal places)


In conclusion, Jeff will need to buy 65.38 more shares to achieve a zero payout policy, bringing his total shares to 865.38 shares.